Commonwealth Construction (CC) needs $1 million of assets to get started, and it expects to have a basic earning power ratio of 30%. CC will own no securities, so all of its income will be operating income. If it so chooses, CC can finance up to 35% of its assets with debt, which will have an 9% interest rate. If it chooses to use debt, the firm will finance using only debt and common equity, so no preferred stock will be used. Assuming a 40% tax rate on all taxable income, what is the difference between CC's expected ROE if it finances these assets with 35% debt versus its expected ROE if it finances these assets entirely with common stock? Round your answer to two decimal places.
Answer #1

Basic Earning Power = EBIT/Total Assets

30% = EBIT/1,000,000

EBIT = $300,000

IF Financed with 35% Debt,

Interest Expenses = 1000,000*35%*9% = $31,500

Income before tax = $268,500

Less: Tax 107,400

Net Income = $161,100

Equity = 650,000

ROE = 24.78%

If financed entirely with equity

Earnings before tax = 300,000

Less: Tax = $120,000

Net Income = $180,000

Equity = 1,000,000

ROE = 18%

Hence, difference in ROE = 24.78% - 18%

= 6.78%